### v2

# A Survey on Employing Large Language Models for Text-to-SQL Tasks

Abstract.
The increasing volume of data stored in relational databases has led to the need for efficient querying and utilization of this data in various sectors. However, writing SQL queries requires specialized knowledge, which poses a challenge for non-professional users trying to access and query databases. Text-to-SQL parsing solves this issue by converting natural language queries into SQL queries, thus making database access more accessible for non-expert users. To take advantage of the recent developments in Large Language Models (LLMs), a range of new methods have emerged, with a primary focus on prompt engineering and fine-tuning. This survey provides a comprehensive overview of LLMs in text-to-SQL tasks, discussing benchmark datasets, prompt engineering, fine-tuning methods, and future research directions. We hope this review will enable readers to gain a broader understanding of the recent advances in this field and offer some insights into its future trajectory.
Large Language Models, Text-to-SQL, Prompt Engineering, Fine-tuning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
In the era of big data, a significant portion of data is stored in relational databases, which serve as the backbone for various organizational data management systems. As the volume of data continues to increase, the capability to efficiently query and leverage this data has emerged as a pivotal factor in enhancing competitiveness across numerous sectors in this era. Relational databases require the use of SQL for querying. However, writing SQL necessitates specialized knowledge, which creates barriers for unprofessional users to query and access databases.
Text-to-SQL parsing is a well-established task in the field of natural language processing (NLP). Its purpose is to convert natural language queries into SQL queries, bridging the gap between non-expert users and database access. To illustrate, imagine a table named cities with three columns: city_name (type: string), population (type: integer), and country (type: string). If we are given the natural language query ”Find all the cities with a population greater than 1 million in the United States,” the Text-to-SQL parsing technique should automatically generate the correct SQL query: ”SELECT city_name FROM cities WHERE population ¿ 1000000 AND country = ’United States’.” Researchers have made substantial progress in this area. Initially, template-based and rule-based methods(Zelle and Mooney, 1996; Li and Jagadish, 2014) were employed. These approaches involved creating SQL templates for various scenarios. While template-based methods showed promise, they required significant manual effort. With the rapid advancement of deep learning, Seq2Seq(Sutskever et al., 2014) methods have emerged as the mainstream approach. Seq2Seq(Sutskever et al., 2014) models provide an end-to-end solution, directly mapping natural language input to SQL output, which eliminates the need for intermediate steps like semantic parsing or rule-based systems. Among Seq2Seq(Sutskever et al., 2014) methods, pre-trained language models(PLMs), which serve as predecessors to large language models (LLMs), show promise in text-to-SQL tasks. Benefiting from the rich linguistic knowledge in large-scale corpora, PLMs become the state-of-the-art (SOTA) solution during that period(Qin et al., 2022).
As model sizes and training data continue to grow, pre-trained language models (PLMs) naturally evolve into large language models (LLMs), exhibiting even greater power. Due to the scaling law and their emergent capabilities, LLMs have made substantial contributions across diverse domains, including chatbots(Dam et al., 2024), software engineering(Hou et al., 2023), and agents(Wang et al., 2024), etc. The remarkable capabilities of LLMs have prompted research into their application for text-to-SQL tasks. Figure  2 shows the evolution of mainstreaming approaches(Lin et al., 2020; Mellah et al., 2021; Rubin and Berant, 2020; Scholak et al., 2021; Qi et al., 2022; Li et al., 2023b; Pourreza and Rafiei, 2024; Dong et al., 2023b; Gao et al., 2023a; Gan et al., 2021c) in the Text-to-SQL domain, represented by the Execution Accuracy in SPIDER test dataset(Yu et al., 2018). Current literature mainly focuses on two main approaches of LLMs: prompt engineering and fine-tuning. Prompt engineering leverages methods like RAG, few-shot learning, and reasoning, while fine-tuning involves training an LLM on task-specific data. There exists a trade-off between prompt engineering and fine-tuning. Usually, prompt engineering demands less data but may lead to suboptimal results, while fine-tuning can enhance performance but necessitates a larger training dataset.
This paper aims to provide a comprehensive survey of the use of LLMs in text-to-SQL tasks. We provide an overview of LLMs’ characteristics and advantages over traditional models, and introduce LLM-based Text2SQL in the following aspects:
  • Text-to-SQL benchmark datasets: We highlight that while previous datasets(Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a) still hold some value in testing the capabilities of LLM, we are in need of more challenging and realistic benchmark datasets. We sum up two notable benchmark datasets, namely BIRD(Li et al., 2024a) and Dr.Spider(Chang et al., 2023), which have been created after the widespread use of LLM and offer new challenges with large scale table schemas and diverse perturbations.
  • Prompt engineering method: We conduct a comprehensive analysis of prompt engineering techniques for text-to-SQL. First , we summarize the basic structure of prompt in text-to-SQL tasks. Subsequently, we explore various approaches employed to improve the effectiveness of prompt engineering methods for text-to-SQL, encompassing the incorporation of diverse supplementary knowledge, the selection of relevant demonstrations, and logical reasoning.
  • Finetuning method: We emphasize the significance of finetuning methods for text-to-SQL and break down finetuning into four main components according to the general fine-tuning process: data preparation, pre-training model selection, model fine-tuning, and model evaluation. It is worth mentioning that in addition to providing a comprehensive analysis of refining methods, we emphasize the potential of utilizing LLMs in both data preparation and model evaluation.
  • Future directions: We discuss the possible future directions of employing LLMs on text-to-SQL tasks, such as privacy concern, autonomous agents, complex schema, benchmarks and domain knowledge.
Our intention is that this survey will serve as a valuable resource for newcomers to this field and offer valuable insights for researchers.
Figure 1. Framework of employing LLMs in Text-to-SQL
2. Overview
2.1. LLMs and LLM-based Text-to-SQL
Large language models(LLMs) have emerged as a milestone for natural language processing and machine learning. The concept of LLMs comes from the practice of continually scaling up the parameter size of pretrained language models (PLMs) and the volume of training data(Zhao et al., 2023), which results in surprising abilities, known as emergent abilities(Wei et al., 2022a), that are not found in smaller PLMs. One example of emergent capability is few-shot learning(Brown et al., 2020), meaning that LLMs can complete the downstream tasks with several proper task demonstrations in the prompt without further training. Another example is the instruction following ability(Wei et al., 2021), with which LLMs have been shown to respond appropriately to instructions describing an unseen task.
Due to the emergent abilities(Wei et al., 2022a) of LLMs and the basic operating principle of LLMs which gradually produce the next word that has the highest probability based on the input prompt(Zhao et al., 2023), prompt engineering becomes one of the two primary streams to apply LLMs to downstream tasks. The representative approaches of prompt engineering are Retrieval Augmented Generation(RAG)(Gao et al., 2023b), few-shot learning(Brown et al., 2020), and reasoning(Wei et al., 2022b; Zhou et al., 2022; Yao et al., 2024). The other stream of LLMs’ application to downstream tasks is called fine-tuning, which follows the ”pre-training and fine-tuning” learning paradigm of PLMs(Zhao et al., 2023) and aims to boost performance in specific domains and address privacy concerns. The general fine-tuning process mainly involves data preparation, pre-training model selection, model fine-tuning, and model evaluation.
Text-to-SQL, which is a challenging task in both the natural language processing(NLP) and database communities and involves mapping natural language questions on the given relational database into SQL queries, has also been revolutionized by the emergence of LLMs. We summarize the general framework of LLM-based Text-to-SQL systems in fig.1. Based on the two main streams of LLM applications, we categorize the methods used in LLM-based Text-to-SQL into two categories, namely prompt engineering and fine-tuning. As for prompt engineering methods, we typically design a well-structured prompt encompassing various components, such as task descriptions, table schemas, questions and additional knowledge, and utilize in-context learning and reasoning methods at the same time. As for fine-tuning methods, we typically generate or collect Text-to-SQL datasets, select appropriate pretrained LLMs and finetuning approaches such as LORA(Hu et al., 2021), and compare the test results before and after to understand the changes in the model’s performance.
2.2. Difference between traditional and LLM-based Text-to-SQL Approaches
Prior to the widespread of LLMs, there were two primary streams for text-to-SQL methods. One stream involves utilizing the sequence-to-sequence (Seq2Seq) model(Sutskever et al., 2014), in which an encoder is designed to capture the semantics of the natural language(NL) question and the corresponding table schema, and a decoder is employed to generate the SQL query token by token based on the encoded question representation. Some notable methods in this approach include IRNet(Jha et al., 2019), SQLNet(Xu et al., 2017), Seq2SQL(Zhong et al., 2017b), HydraNet(Lyu et al., 2020), Ryansql(Choi et al., 2021), Resdsql(Li et al., 2023b) and ISESL-SQL(Liu et al., 2022a). The other stream involves finetuning PLMs such as BERT(Devlin, 2018), which leverage the extensive knowledge present in large-scale text collections and have proved to be effective in enhancing the performance of downstream text-to-SQL parsing tasks.
We argue that there are two distinct aspects that differentiate traditional text-to-SQL approaches from LLM-based ones:
  • Novel Paradigm: Traditional text-to-SQL approaches necessitate training, whereas LLMs can often circumvent this requirement. Leveraging the instruction following ability(Wei et al., 2021) of LLMs, LLMs can accomplish the text-to-SQL tasks with the aid of appropriate instructions and information.
  • Uniform Architecture: According to a previous survey(Qin et al., 2022), the encoder and decoder in traditional approaches can be designed using diverse architectures such as LSTM(Hochreiter and Schmidhuber, 1997), Transformer(Vaswani et al., 2017), and even GNNs(Bogin et al., 2019). In contrast, LLMs adhere to a uniform transformer-based architecture, which not only enables easier scaling up but also facilitates a more streamlined implementation.
2.3. Why LLM-based Text-to-SQL
There has been a sharp increase in the number of LLMs employed for text-to-SQL tasks in recent times. After conducting a comprehensive survey of recent papers, we identify several key reasons for this trend, which are summarized below:
  • Enhanced Performance: Figure  2 illustrates the progression of mainstreaming approaches within the Text-to-SQL domain, as indicated by the Execution Accuracy in the SPIDER test dataset(Yu et al., 2018). As we can see, LLM-based methods have significantly improved the SOTA performance, demonstrating the remarkable capabilities of LLM-based methods. The listed references(Lin et al., 2020; Mellah et al., 2021; Rubin and Berant, 2020; Scholak et al., 2021; Qi et al., 2022; Li et al., 2023b; Pourreza and Rafiei, 2024; Dong et al., 2023b; Gao et al., 2023a; Gan et al., 2021c) support this observation.
  • Generalization Ability and Adaptability: As we mentioned above, LLMs have introduced a new paradigm, namely prompt engineering, which benefits from their ability to follow instructions(Wei et al., 2021), making LLMs easily transferable to different settings without additional training. Furthermore, the in-context learning ability(Zhao et al., 2023) of LLMs further enhances their generalization and adaptability capabilities, as they can learn from examples provided to seamlessly fit into various settings.
  • Future Improvements: LLM-based methods hold a promising future for advancements. Since the global community is prioritizing the enhancement of LLMs, efforts and resources are concentrated towards supporting research on LLMs. This includes scaling up LLMs, creating new prompting methods, generating high-quality and extensive datasets, and fine-tuning LLMs across various tasks. The progress made by the LLM community will undoubtedly and continuously drive LLM-based text-to-SQL methods to new SOTA.
Figure 2. The evolution of Text-to-SQL approach over time
3. Benchmark
High-quality datasets play a crucial role in the training and testing of Text-to-SQL tasks. Previous datasets like (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a) made pivotal contributions to the advancement of Text-to-SQL domain research, particularly before the emergence of ChatGPT and GPT-4(Achiam et al., 2023). However, with the introduction of ChatGPT and GPT-4(Achiam et al., 2023), LLM began to be widely used for Text-to-SQL tasks, thereby continuously elevating the state of the art (SOTA) for such tasks. For instance, the SOTA of the Spider(Yu et al., 2018) dataset (accuracy of execution results) increased from approximately 73 to 91.2. Even though datasets from (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a) still hold some value for testing LLM’s capabilities, we require more challenging, realistic benchmark datasets. Here, we’ve summed up two noteworthy benchmark datasets born after the widespread use of LLM.
3.1. BIRD
Most datasets (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a) focus on the database schema with few rows of database values, leaving the gap between academic research and real-world applications. To address this gap, (Li et al., 2024a) introduced the BIRD dataset. The BIRD dataset includes 12,751 text-to-SQL pairs and 95 databases, with a total size of 33.4 GB across 37 professional domains. It also highlights new challenges including dirty and noisy database values, external knowledge grounding between natural language (NL) questions and database values, and SQL efficiency, especially in the context of massive databases. Experimental results indicate that even the most effective text-to-SQL models, i.e. GPT-4(Achiam et al., 2023), only achieve 54.89% in execution accuracy, which is still far from the human result of 92.96%, proving that challenges still stand.
3.2. Dr.Spider
The work in (Chang et al., 2023) pointed out that Text-to-SQL models are vulnerable to task-specific perturbations, but previous curated robustness test sets usually focus on individual phenomena. To bridge this gap, (Chang et al., 2023) introduced the Dr.Spider dataset. Based on the Spider dataset(Yu et al., 2018), Dr.Spider designs 17 perturbations on databases, natural language questions, and SQL queries to measure robustness from various angles. To collect more diversified natural question perturbations, (Chang et al., 2023) leveraged large pre-trained language models (PLMs) to simulate human behaviors in creating natural questions in a few-shot manner. Experimental results indicate that even the most robust model encounters a 14.0% overall performance drop and a 50.7% performance drop on the most challenging perturbation.
4. Prompt Engineering
Prompt engineering, sometimes called in-context learning, means changing LLM’s output on certain tasks by changing prompt words when interacting with LLM. Due to the autoregressive decoding properties(Fu et al., 2024) of the vast majority of LLMs, they predict the following text based on all currently visible preceding text which is also called context. We can give the autoregressive decoding formula 1, as $y_{t}$ indicates the next token LLM will output, and $x$ indicates prompt tokens given by users. The design of prompt words as context will affect the probability distribution of all upcoming tokens, thereby affecting the final generation.
(1) $y_{t}=\arg\max P(y_{t}|y_{1:t-1},\textbf{x})$
Most of the better prompt words come from the experience of practitioners and the detailed description of the task itself. Prompt engineering costs much less compared to finetuning methods, has strong generalization potential for universal Text2SQL tasks, and obtains considerable accuracy in SQL generation.
So far, many works have applied prompt engineering to Text2SQL scenarios. These works focus on how to integrate natural language and related SQL experience with problems into LLM. The specific application directions are shown in Fig 3 and analyzed as follows.
Figure 3. Taxonomy of Prompt Engineering in Text-to-SQL
4.1. Basic Structure of Prompt Engineering in Text2SQL
The input for all text generation tasks has some essential components, such as providing a story title or beginning for story continuation, and providing a problem description and programming language for code generation. The essential elements in Text2SQL tasks include two types: the natural language problem itself and the organizational form (schema) of the database. Recent works(Arora et al., [n. d.]; Rajkumar et al., 2022) introduced the prompt representations for the five most commonly used forms of database organization.
Question Some works (Hu et al., 2023; Rajkumar et al., 2022) only included prompts for natural language problems themselves and not any other information from the database. It will also be designed as a type of prompt, with a SELECT ending to indicate that the SQL output by LLM does not include special symbols, which is usually used for comparative experiments and is less effective compared to other prompt words.
API Docs Some works (Guo et al., 2023a; Liu et al., 2023a; Rajkumar et al., 2022; Tai et al., 2023; Wang et al., 2023a) used the prompt template of Text2SQL task provided in the Codex official API documentation, with each line starting with the comment symbol of SQLite, followed by the form of Table (Column1, Column2,…). After the table description is completed, the last row provides the SELECT word.
SELECT X The work in (Rajkumar et al., 2022) represented the prompt as a ”SELECT * From Table LIMIT X” statement and the running result of this SQL statement, where X is an integer and is often taken as 3 based on experience. The purpose is to enable LLM to understand some of the data in the database and conform to the format of the data itself when generating SQL.
CREATE TABLE The works (Li et al., 2024a; Nan et al., 2023; Rajkumar et al., 2022; Gao et al., 2023a) represented prompt words for database organization as CREATE TABLE creation statements. Explicitly including the data type of each column and possible primary and foreign key relationships in the statements will align the data format and perform multi-table connections when LLM generates SQL. Other works (Cheng et al., 2022; Tai et al., 2023; Xie et al., 2024) used CREATE TABLE statements and partial data to build prompt words, which is more in line with the real situation of data engineers writing SQL.
Others In addition to the four commonly used representations mentioned above, the work in (Arora et al., [n. d.]) concatenated the prompt words of API Documents and SELECT X as its own prompt words; (Arora et al., 2023) added random sampling of categorical columns and range information for numerical columns after CREATE TABLE statements; Some works (Zhang et al., 2023; Dong et al., 2023b; Nan et al., 2023; Sun et al., 2023; Xie et al., 2024) separated primary or foreign keys from database descriptions and described them separately, attempting to increase LLM’s attention to primary and foreign keys; The work of (Kothyari et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b; Guo et al., 2023b; Sun et al., 2023) uses the form of Table.Column to list database schema, which is usually used in conjunction with other supplementary knowledge filling methods about schema; (Chang and Fosler-Lussier, 2023) modified the form of SELECT X to list column-wise data, for the purpose of explicitly enumerate categorical value.
4.2. Supplementary Knowledge Prompt
Despite being equipped with the requirements of question and database, LLM cannot achieve advantages over traditional deep models in the Text2SQL task. To improve the effectiveness and potential of LLM, many works have begun to add supplementary and professional knowledge for the Text2SQL task in prompt words, so that the SQL generated by universal LLM will be more accurate and reliable. In fact, this is also the key to the prompt engineering method. Some of the knowledge is reflected in adding domain-related rules, experiences, and restrictions, while others are about removing the useless parts of the requirements based on specific problems, making LLM more focused on the useful parts. Specifically, it can be divided into the following categories.
4.2.1. Schema Knowledge
Whether researchers are predicting research datasets in models, or data engineers are writing SQL in real-world scenarios to solve Text2SQL tasks, they all face the challenge of having too many databases, too long schema, and difficulty in identifying effective table and column information. The major approach to solving this problem is to focus the model’s attention on schema related to the problem and ignore irrelevant schema, known as the schema linking method. This method is considered to incorporate schema knowledge into the prompt words. Many works (Zhang et al., 2023; Dong et al., 2023b; Kothyari et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Nan et al., 2023; Tai et al., 2023; Wang et al., 2023b; Sui et al., 2023; Li et al., 2024b) proposed their customized approach on schema linking. Research(Zhang et al., 2023) represents all schema items in the form of Table.Column, and then calculates their similarity to the current problem separately. The tables and columns with the highest similarity are added to the prompt; At the same time, the similarity score between each table and the current problem is also calculated, and the table with the highest score is added to the prompt. These most relevant schema items will guide LLM in the Chain-of-Thought process of generating SQL when solving new problems. (Dong et al., 2023b) built three key modules on the Text2SQL task based on ChatGPT, and used a two-step schema linking in one of its modules. Specifically, the first step is to use ChatGPT to recall the tables most relevant to the issue, and the second step is to use ChatGPT to recall the columns most relevant to the question. CRUSH4SQL (Kothyari et al., 2023) first requires LLM to use only the original natural language problem as the prompt, add relevant examples, and generate a minimal database schema that can be used to solve the problem in a ”hallucinatory” manner. Then, using this ”hallucinatory” schema as an index, similarity retrieval is performed on the real schema to obtain subsets related to the problem in the real schema. DIN-SQL(Pourreza and Rafiei, 2024) utilizes the SPIDER dataset to generate examples as prompt words in the schema-linking scenario. The author organizes the examples in the form of Chain-of-Thought(Wei et al., 2022b), including natural language questions in SPIDER, keywords extracted from the questions, and table schema corresponding to the keywords. (Nan et al., 2023) has enhanced the semantics of the schema by representing the relationships between tables in the form of Entity Relationship, and specifically explaining which tables can be joined. (Tai et al., 2023) first breaks down Text2SQL into several sub-problems, and generates the corresponding most likely schema for each sub-problem. (Sui et al., 2023) adopts the Explain-Squeeze method, The Explain section uses LLM to answer schema-related questions step by step, including information dimensions in the query, identifying tables in the query, listing and explaining the columns in these tables, and locating the columns that appear in the query; The Squeeze section extracts the final answer from LLM’s answers to the above questions. (Li et al., 2024b) perform schema linking after generating the PreSQL statement in the first step, and then use the selected schema for prompt simplification, and finally use the simplified prompt to generate SQL for LLM.
4.2.2. SQL Knowledge
The ultimate goal of the Text2SQL task is to generate SQL statements that are grammatically correct, have correct results, and run efficiently. To enhance the understanding of SQL syntax and writing rules for LLM, many works (Arora et al., [n. d.], 2023; Zhang et al., 2024; Dong et al., 2023b; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Nan et al., 2023; Gu et al., 2023; Xie et al., 2024) have added SQL-related rules and limitations to prompt words, collectively referred to as SQL knowledge. The approach proposed in (Arora et al., 2023, [n. d.]) uses SQL sets to cover the complete SQL clauses, operators, and functions. It uses the smallest set that can cover the data domain to the maximum extent to put into the prompt. Finally, it is combined with the test query to form a ”generic prompt”, and the format adopts the SQLite lattice. (Arora et al., [n. d.]) also adds rules such as removing excess spaces to prevent interference with SQL running and filtering condition judgment; (Arora et al., 2023) adapts the existing SQL set to the schema of the test data to generate SQL corresponding to the structural position and using natural language to annotate them. The resulting ”text-SQL” example is then decomposed into natural language using the Least-to-Most(Zhou et al., 2022) method and represented in the middle using NatSQL(Gan et al., 2021c) to obtain a multi-step decomposed Text2SQL example. A new decoding algorithm is proposed by (Zhang et al., 2024) to address the ambiguity issue in SQL. Multiple SQL templates with diversity are designed on the decoding side, and multiple valid answers to the original question are obtained through beam search. When generating SQL, several rules can be added to optimize the quality of the generated SQL, including reducing redundant columns in SELECT, reducing redundant INTERSECT execution results, disabling Left JOIN, and switching to JOIN(Dong et al., 2023b). (Pourreza and Rafiei, 2024; Xie et al., 2024) use JOIN and subqueries as standards, categorizing each problem and using corresponding examples to guide LLM in generating SQL for each type of problem. (Nan et al., 2023) also adopts classifying according to difficulty, using the classification criteria of the dataset SPIDER(Yu et al., 2018), and using the kmeans(Ahmed et al., 2020) method to select examples in each classification. (Liu and Tan, 2023) proposed a method of gradually generating SQL prompt words in the order of clauses. A divide-and-conquer framework for Text2SQL was proposed by (Gu et al., 2023), which divides the Text2SQL task into a Structure stage and a Content stage. The Structure stage allows the language model to generate an SQL framework containing SQL keywords, operators, and placeholders; The Content stage allows the language model to generate specific numerical values for these placeholders, including column names, table names, and constants.
4.2.3. Other Language Knowledge
The work (Cheng et al., 2022) takes a different approach and chooses to have LLM generate the new language BINDER-SQL designed for the original problem first. BINDER-SQL is consistent with the SQL framework. However, the difference is that some column names and values are replaced with API expressions, which correspond to sub-problems and some information in the original table. LLM is used to predict the answers to sub-problems and information, and the results are filled into BINDER-SQL, ultimately converted into executable SQL statements. The author designed specialized lexical analysis, syntax analysis, and executors for the BINDER-SQL language, and used the few-shot method(Brown et al., 2020) to generate BINDER-SQL for LLM.
4.2.4. Task-related Knowledge
Some external task-related knowledge will also be helpful when generating domain-specific SQL statements. For example, in the scenario where customers purchase goods, (Li et al., 2024a) adds a calculation method for a specific concept of ”price per unit of product” to the prompt words which helps generate more accurate SQL when answering questions about that concept.
4.2.5. Memory Knowledge
ChatDB (Hu et al., 2023) uses symbolic memory to assist LLM in complex multi-step reasoning. Specifically, it decomposes a complex database operation task into multi-step execution and saves the results of SQL execution in memory, which determines the next step of LLM’s database operation.
4.2.6. Question Knowledge
Some work (Guo et al., 2023a, b; Sui et al., 2023; Gao et al., 2023a; Xie et al., 2024; Li et al., 2024b) start from the question itself, hoping to gain experience in interacting with similar questions through problem mining, to guide LLM to achieve better performance on current problems. This type of method is usually used in conjunction with few-shot, and the experience of similar questions will often be reflected in prompt words in the form of examples. (Guo et al., 2023a, b) propose a GPT-3.5 framework based on examples and reasoning to precisely control the example-related and example-independent knowledge of Text2SQL. The design of specific prompt words is based on desemantizing the input problem to obtain the question intention, and then using KNN(Guo et al., 2003) to adaptively retrieve examples related to the problem intention, to ensure the relevance of the retrieved examples to the current problem and sufficient information after adding prompt words. (Li et al., 2024b) also utilizes the skeleton of questions to help retrieve similar ”text-SQL” pairs. (Sui et al., 2023) adopts the method of query rewriting to rewrite the question to a more suitable understanding for LLM. It is divided into three steps: first, extract the fuzzy vocabulary in the problem, then use the context or internal knowledge of LLM to rewrite the above vocabulary as explicitly defined terms, and finally replace the words uniformly. The results of each step are obtained independently through interaction with LLM.
4.3. Example Selection
If we want LLM pre-trained on general domain data to achieve good results in the Text2SQL domain, providing examples for the Text2SQL domain is a good solution, just like humans are good at imitating similar situations to make more accurate decisions. The LLM-related work in the field of Text2SQL is divided into two categories in the selection of examples: One kind of work (Liu et al., 2023a; Dong et al., 2023b; Hu et al., 2023; Wang et al., 2023a; Liu and Tan, 2023; Rajkumar et al., 2022; Chang and Fosler-Lussier, 2023; Sui et al., 2023) adopts zero-shot strategy, which does not include Text2SQL examples in prompt words, but focuses on other aspects like the table column structure of the database and dividing Text2SQL problems into multiple stages to solve; The other kind of work (Guo et al., 2023a; Arora et al., [n. d.]; Zhang et al., 2023; Arora et al., 2023; Cheng et al., 2022; Li et al., 2024a; Kothyari et al., 2023; Pourreza and Rafiei, 2024; Nan et al., 2023; Tai et al., 2023; Gu et al., 2023; Wang et al., 2023b; Sun et al., 2023; Gao et al., 2023a; Xie et al., 2024; Li et al., 2024b) adopts few-shot strategy(Brown et al., 2020), which includes several Text2SQL examples in the design of prompt words. It is hoped that LLM can learn the correct pattern for generating SQL from the examples, which can be adopted in many subdivision stages of Text2SQL. Since zero-shot methods are usually only used to save the number of tokens consumed by LLM and reduce costs, the following will mainly introduce the existing work on the application of few-shot methods.
Few-shot The few-shot method can be used to improve the effectiveness of nearly every corresponding step in the Text2SQL task. (Arora et al., [n. d.]; Li et al., 2024a; Sun et al., 2023) select the combination example of database schema, text query, and SQL statements which provides direct guidance for Text2SQL tasks. To reduce the number of tokens input into LLM and save costs, sometimes the database schema was removed in the examples (Nan et al., 2023; Gao et al., 2023a; Xie et al., 2024), while text query and SQL statements were retained. (Guo et al., 2023a, b) only select a few SQL statements and add prompt words to allow LLM to have a comprehensive understanding of SQL syntax, keywords, and possible solutions to the current problem. (Zhang et al., 2023; Arora et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b) merge the text query, SQL statements, and inference processes into the example to guide LLM in inferring the current problem according to the given pattern. (Cheng et al., 2022) apply the few-shot method to the generation of its own BINDER-SQL, allowing LLM which is completely unfamiliar with the BINDER-SQL language to understand the common writing methods of the language through examples, laying the foundation for its subsequent API calls and SQL generation. Unlike providing examples related to SQL, (Kothyari et al., 2023) focuses on creating illusions for LLM when inferring schemas, with examples in prompt words including text queries and their corresponding nonreal schemas. Similar to using the few-shot method in schema prediction, (Pourreza and Rafiei, 2024) used partial text queries and corresponding real schema from the SPIDER dataset to form the examples in the prompt. (Gu et al., 2023) using prompt learning, divide SQL generation into two parts: SQL structure and SQL content, and use feed shot to solve them separately.
4.4. Reasoning
Similar to the application of LLM in other scenarios, The application of LLM in the Text2SQL field can adopt a multi-step reasoning pattern, generating answers that gradually approach the correct SQL step by step. The main multi-step reasoning methods include Chain-of-Thought (CoT)(Wei et al., 2022b), Least-to-Most(Zhou et al., 2022), Self-Consistency(Wang et al., 2022b) and Self-Correction.
4.4.1. Chain of Thought (CoT)
Chain-of-Thought(CoT) guides LLM to think step by step and generate answers by adding statements such as ”Let’s think step by step” to the prompt words, allowing LLM to execute multiple steps in sequence and generate the final result in the last step. A lot of work (Zhang et al., 2023; Li et al., 2024a; Hu et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Wang et al., 2023b; Xie et al., 2024) adopted the CoT method to help LLM generate SQL. ACT-SQL (Zhang et al., 2023) utilizes LLM’s few-shot method to automatically extend ordinary examples into CoT pattern examples. Through multiple rounds of LLM interaction, the SQL generated from the previous interaction will be executed and affect the in-memory database(Hu et al., 2023). The next interaction will use the current in-memory database status as prompt words, allowing LLM to decide the next database operation. DBCopilot (Wang et al., 2023a) guides LLM to select appropriate tables and columns before writing SQL in the example, allowing LLM to generate two-step CoT answers. In (Liu and Tan, 2023), three types of CoT solutions were proposed for LLM to solve the Text2SQL problem, namely clause-by-clause generation, selecting tables and columns before generation, and refining after completion.
4.4.2. Least-to-Most
Least-to-Most breaks down a complex problem into a series of simpler sub-problems and then solves these problems to achieve the effect of solving the original complex problem. (Arora et al., 2023) uses LLM to automatically extend existing examples to generate Least-to-Most pattern examples. (Tai et al., 2023) divides the question based on the connecting words and prepositions in the original question so that LLM can generate or add a clause based on the previous answer when answering the sub-question, and finally concatenate it into SQL that meets the grammar requirements.
4.4.3. Self-Consistency
The Self-Consistency method mainly adopts a majority voting strategy, allowing LLM to generate multiple answers to the same question and select the answer that appears the most frequently as the final answer. In (Dong et al., 2023b; Sun et al., 2023), the Self-Consistency method was directly used to perform a major vote on the SQL generated by its LLM, achieving good performance improvement. PET-SQL (Li et al., 2024b) proposes a cross-consistency method to instruct several LLMs under lower temperatures to generate SQLs and then make votes across the SQLs’ executed results.
4.4.4. Self-Correction
After LLM generates an answer, the Self-Correction method uses rules under specific questions and tasks to have LLM check the correctness of the answer; In the Text2SQL scenario, the Self-Correction method can use SQL-related rules for checking, and can also provide LLM with the results or error logs generated by running SQL statements for inspection. The approach of (Guo et al., 2023a, b) is to modify the small number of schema in the original prompt words to include a full number of schema in the prompt words and generate SQL again if the SQL generated by LLM cannot run. In (Pourreza and Rafiei, 2024; Wang et al., 2023b), the incorrect SQL and the information indicating the SQL error will be used as prompts for LLM to regenerate the SQL.
5. Fine tuning
Although large language models (LLMs) like GPT-4(Achiam et al., 2023) have achieved significant success in Text-to-SQL tasks using prompting methods such as RAG, ICL, and CoT, it’s important to note that there is a certain gap between these low-cost prompt methods and fine-tuning LLMs(Sun et al., 2023; Meng et al., 2023). At the same time, concerns about privacy issues often arise with API call models(Xue et al., 2023). Therefore, we should also pay attention to the application of fine-tuned LLMs in Text-to-SQL tasks. (Sun et al., 2023) fine-tuned the PaLM-2(Anil et al., 2023) model for Text-to-SQL tasks, achieving unprecedented performance for the Spider dataset at the time; (Wang et al., 2023b) constructed an agents framework for Text-to-SQL tasks and fine-tuned the code-llama(Roziere et al., 2023) based on this; and (Gao et al., 2023a) fine-tuned several open-source LLMs, demonstrating the huge potential of fine-tuned LLMs for Text-to-SQL tasks. We will organize this section according to the general fine-tuning process, including data preparation, pre-training model selection, model fine-tuning, and model evaluation and fig.4 shows a taxonomy tree that summarizes the structure and contents of this section.
Figure 4. Taxonomy of Fine-tuning in Text-to-SQL
5.1. Data Preparation
The quality of the training data plays a significant role in determining the upper limit of the fine-tuning model’s effectiveness, making ”data preparation” a crucial step in the entire fine-tuning process. The fine-tuning dataset can be obtained by integrating existing datasets or creating new ones.
5.1.1. Integrating Existing Datasets
When integrating existing datasets, our goal is usually to acquire a large-scale, general-purpose model in the text2sql domain, i.e., an LLM with strong cross-domain text2sql capabilities. When training a general-purpose text2sql model, we can use common existing datasets (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a; Chang et al., 2023) for fine-tuning. To improve the model’s generalization, different datasets have adopted various data enhancement methods. For example, Spider-SYN(Gan et al., 2021b) and Spider-Realistic(Deng et al., 2020) can eliminate cases where the database schema explicitly appears in the question, Spider-DK(Gan et al., 2021a) and ADVETA(Pi et al., 2022) require additional knowledge or have undergone adversarial modifications to the database column names, Spider-CG(Gan et al., 2022) focus on improving the model’s combinatorial generalization ability, SParC(Yu et al., 2019b) and CoSQL(Yu et al., 2019a) include multi-turn dialogue text2sql tasks, CSpider(Min et al., 2019) and DuSQL(Wang et al., 2020) enhance the model’s multilingual capabilities, Dr.Spider(Chang et al., 2023) have made multi-faceted adversarial modifications to the question, column names, and SQL, and Bird(Li et al., 2024a) simulate real text2SQL scenarios with multiple tables, complex questions, and noise. It’s worth noting that (Jang et al., 2023) pointed out that due to negative task transfer, training on more tasks may not result in better generalization than training on a single task. Based on this conclusion, we believe that training on multiple datasets may encounter similar problems. Also, due to the significant improvement of CoT on inference effects, the fine-tuning dataset needs to contain a certain amount of CoT. (Chung et al., 2024) pointed out that a certain amount of CoT data in the instruction fine-tuning dataset is enough to improve the model’s reasoning ability, but all CoT data will decrease the non-reasoning inference ability.
5.1.2. Building New Datasets
When building new datasets, the dataset can be used to train both a general-purpose model in the text2SQL domain and a specialized LLM for a specific domain. Traditional methods of building new datasets often involve manual works, such as Spider(Yu et al., 2018), which was manually annotated by 11 Yale University students. Through this method of humans posing questions and annotating the correct SQL, we have obtained high-quality datasets like (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a). However, manually annotating datasets is inefficient and costly (Xu et al., 2023b). With the advancement of NLP technology, especially the emergence of LLM, a batch of semi-automatic and fully automatic dataset construction methods have been born. (Chang et al., 2023) builds a new dataset based on Spider by first using crowdsourcing to disturb the dataset, then manually summarizing the types of disturbances based on the new data obtained, and finally using the LLM to generate new disturbances in a few-shot manner under each type. (Hu et al., 2024) investigated a fully automatic approach to create datasets for text2SQL tasks. The method followed a paradigm that involved providing schema and example rows of databases to LLMs, generating SQL queries using zero-shot synthesis by LLMs, and validating the queries through execution-based validation.
As of the time of our writing, we have not found other automatic methods for constructing datasets for text2sql tasks. Therefore, we have summarized several methods of generating instruction fine-tuning datasets that may be transferable to text2sql tasks, which may benefit future work. Existing automatic methods of generating instruction fine-tuning datasets(Wang et al., 2022a; Honovich et al., 2022; Xu et al., 2023a) generally adopt the pattern of ”sampling from existing data, constructing more data in a few-shot manner, and filtering”. One point worth noting is that (Xu et al., 2023b) pointed out that due to the proportion of experts and the time and effort involved in constructing data, humans may find it difficult to compile high-complexity data. (Xu et al., 2023b) constructs data by increasing instructions vertically to increase the difficulty of the problem and expanding the problem horizontally to increase the richness of the problem, thereby generating a complex and rich instruction dataset step by step.
5.2. Selection of Pretrained Models
Numerous open-source language models are available, such as (Touvron et al., 2023; Roziere et al., 2023; Li et al., 2023a; Nijkamp et al., 2022; Guo et al., 2024; Anil et al., 2023). These models vary in terms of parameter scales, pre-training corpora, and pre-training processes, and all can be fine-tuned. For example, the PaLM-2(Anil et al., 2023) model was fine-tuned for the text2sql task as detailed in (Sun et al., 2023), and it achieved the best results on the Spider dataset at that time. The code-llama(Roziere et al., 2023) model was fine-tuned in (Wang et al., 2023b) and was used in the Agents system, which was designed for the text2sql task. However, when selecting a model for fine-tuning, it is crucial to consider a variety of factors, including the pre-trained model’s capabilities, the volume of fine-tuning data, the model parameters’ scale, and computational capabilities. As mentioned in (Chung et al., 2024), the advantages of fine-tuning may be attributed to its capacity to more effectively utilize the knowledge gained during pre-training. Consequently, selecting a language model with a more extensive pre-training corpus could be advantageous. (Dong et al., 2023a) indicates that pre-trained models with larger parameter scales perform better when given the same amount of training data. Research in (Zhang et al., 2024) found that coding-specific models are more effective in Text-to-SQL tasks.
5.3. Model Training
In the field of text-to-SQL, model fine-tuning can adopt general fine-tuning methods, mainly including Fully Fine-tuning(FFT) and Parameter-Efficient Fine-tuning(PEFT)(Liu et al., 2023b, 2021; Hu et al., 2021; Dettmers et al., 2024; Liu et al., 2022b; Lester et al., 2021). For example, (Wang et al., 2023b) and (Sun et al., 2023) tuned LLMs using FFT while (Zhou et al., 2024) tuned LLMs using LoRA(Hu et al., 2021) and QLoRA(Dettmers et al., 2024). PEFT aims to improve training efficiency and reduce training costs by only fine-tuning a small number of model parameters instead of all the model’s parameters. According to research in (Jang et al., 2023), PEFT is less prone to catastrophic forgetting compared to FFT during use, thus proving its superiority.
5.4. Model Evaluation
After fine-tuning the model, an important step is to compare the before and after results to understand the changes in the model’s performance. In the context of text2sql tasks, the common approach is to conduct a comprehensive metric analysis on the test set, such as calculating the accuracy rates of EM and EA before and after fine-tuning. However, beyond these metrics, detailed analysis can reveal more in-depth insights. For instance, we can explore the performance of LLM in various scenarios by categorizing the inputs. The emergence of LLM also opens up new possibilities for model evaluation analysis. Therefore, we categorize the methods of model evaluation into three main types: metric analysis evaluation, category analysis evaluation, and LLM-based analysis evaluation.
5.4.1. Metric Analysis Evaluation
We have summarized the following four approach metrics:
Exact Set Match Accuracy (EM) is determined by comparing the literal content of the generated SQL and the ground-truth SQL. Specifically, EM compares whether the SQL clauses of the generated SQL and the ground-truth SQL are consistent. However, there are multiple ways to express the same SQL problem, so the EM metric often underestimates the prediction accuracy of the model.
Execution Accuracy (EX) is concluded by comparing the execution results of the generated SQL and the ground-truth SQL. However, SQL with different logic may yield the same results when executed, so EX may also overestimate the prediction accuracy of the model.
Test-suite Accuracy(TS)(Zhong et al., 2020) creates a small, concentrated database test suite from a large number of randomly generated databases. These databases have a high code coverage rate for accurate queries. During the evaluation process, it measures the annotation accuracy of the predicted queries in this test suite, effectively calculating the strict upper limit of semantic accuracy.
Valid Efficiency Score(VES)(Li et al., 2024a) includes SQL execution efficiency in the evaluation scope. The calculation formula is shown in equation 2, where the hat symbol represents the predicted result, 1 is the indicator function, which is only 1 when the predicted SQL is equivalent to the correct SQL, and R is the square root of the ratio. Intuitively, the higher the correctness rate of the generated SQL, the higher the execution efficiency of the generated SQL, and the higher the VES value.
(2) $VES=\frac{\sum^{N}_{n=1}\mathbf{1}(V_{n},\hat{V}_{n})\cdot\mathbf{R}(Y_{n},\hat{Y}_{n})}{N},\mathbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{\frac{E(Y_{n})}{E(\hat{Y}_{n})}}$
ESM+(Ascoli et al., 2024) is proposed based on Exact Set Match Accuracy (EM) and applies new rules to LEFT JOIN, RIGHT JOIN, OUTER JOIN, INNER JOIN, JOIN, DISTINCT, LIMIT, IN, foreign keys, schema check and alias checks compared to EM. (Ascoli et al., 2024) compare the performance of 9 LLM-based models using TS, EM, and ESM+ and the results indicate that ESM+ offers a substantial improvement by reducing the occurrences of both false positives and false negatives that commonly plague the earlier metrics(EM and TS).
5.4.2. Category Analysis Evaluation
In addition to these above criteria, granular analysis may also be beneficial. Through categorical analysis, we can discern the varying capabilities of the model from different perspectives.
Classification by Dataset Characteristics.(Liu et al., 2023a) considered the characteristics of datasets (Yu et al., 2018; Gan et al., 2021b, a; Deng et al., 2020; Gan et al., 2022; Pi et al., 2022; Min et al., 2019; Wang et al., 2020; Yu et al., 2019b, a; Zhong et al., 2017a), and separately tested (1) common zero-shot scenarios, (2) scenarios that eliminate the explicit appearance of the database schema in the questions, (3) scenarios that require additional knowledge or adversarial modifications applied to the database column names, (4) the compositional generalization scenario, (5) multi-turn dialogue scenarios, and (6) multi-language scenarios.
Classification by Prompt Engineering Methods.(Gao et al., 2023a) tested the effects of text2sql under different prompt formats, different example selection methods, and different example organization methods.
Classification by Question Difficulty.Spider (Yu et al., 2018) proposed a set of SQL difficulty classification standards, which consider SQL with more keywords, nested subqueries, and filter conditions to be more difficult. Based on these standards, the questions in the Spider dataset were divided into easy, medium, hard, and extra hard difficulties.
Classification by Error Type.Unlike the above classification methods, (Zhang et al., 2024) did not classify the data, but classified the error types of the output results. (Zhang et al., 2024) divided the errors into System Error and Result Error. The former refers to errors discovered by the database engine during the SQL execution process, while the latter refers to errors in the execution results. (Zhang et al., 2024) further subdivided execution result errors into table query errors, column selection errors, join column errors, and filter condition errors.
5.4.3. LLM-based Analysis Evaluation
(Chiang and Lee, 2023) found that due to its powerful natural language processing capabilities, LLM has the potential to evaluate models. LLM Comparator(Kahng et al., 2024) proposed a system that can visualize and evaluate the differences between two models, analyzing the differences from the ”when, why, how” dimensions. Although the method of LLM Comparator(Kahng et al., 2024) is not specifically for the text2sql task, we can apply the above problem classification methods (Liu et al., 2023a; Gao et al., 2023a; Yu et al., 2018; Zhang et al., 2024) and macro standards design (Yu et al., 2018; Zhong et al., 2020; Li et al., 2024a) to the evaluation system of LLM Comparator(Kahng et al., 2024), thereby migrating the evaluation system to the text2sql field.
6. Future Direction
Though the application of Large Language Model(LLM) in Text-to-SQL tasks has seen impressive advancements, developing high-quality text-to-SQL parsers continues to present multiple complexities. In this section, we intend to delve into potential future directions and share our perspectives.
6.1. Privacy Concern
ChatGPT and GPT-4(Achiam et al., 2023) demonstrate impressive capabilities, acknowledged as the most powerful Large Language Models (LLMs) available today. However, concerns over privacy issues often arise while using API call models(Xue et al., 2023). When executing Text-to-SQL LLM applications in the industrial setting, it may require the transfer of KPIs, table schemas, sample data, even specific knowledge to the LLM API provider, such as OpenAI. Undoubtedly, this could lead to potential leakage of trade secrets. Private deployment of LLM can serve as a solution for this problem. However, there remains a gap between the capabilities of open-source models and ChatGPT or GPT-4(Achiam et al., 2023). In the presence of adequate data, fine-tuning an LLM emerges as a feasible option, but it could potentially lead to a compromise in the model’s other capabilities or even result in catastrophic forgetfulness.
6.2. Autonomous Agents
LLM, under the training of large-scale corpora, has emerged with powerful generalization capabilities. With the help of the ReAct(Yao et al., 2022) framework, we can build LLM-powered Autonomous Agents. These agents can automatically complete tasks, a process that can be modeled as a state transition path. LLM, acting as the brain, controls the entire process by interacting with the environment, humans, or other agents, while referencing past information stored in Memory to determine the current state and decide which predefined tool to use to achieve state transition. Equipped with self-correction, LLM-powered Autonomous Agents can check the validity and rationality of past behaviors and explore other possible decisions when a single-step decision fails. Compared to pipeline systems, LLM-powered Autonomous Agents have higher flexibility and explorability.
Why does Text-to-SQL need the flexibility and explorability of Autonomous Agents? We examined the process of humans writing SQL and compared it with the existing Text-to-SQL process. Firstly, when humans write SQL, they write some simple SQL to understand what the data looks like according to the task needs, while the existing Text-to-SQL process often ignores this step. Works like (Cheng et al., 2022; Rajkumar et al., 2022) obtain some data from the database through ”SELECT * FROM Table LIMIT X”, but this fixed method is not close to the task itself. Humans can adjust SQL to observe more interesting data as needed. Secondly, when humans write SQL, it’s not done in one step. Humans will try to write SQL, execute SQL to get results, observe the results to confirm the correctness of SQL, modify SQL, and repeat the above steps until the requirements are met. However, the existing Text-to-SQL methods lack observation of the results and autonomous judgment of the correctness of the results. However, as of the time of writing, we have not found the application of LLM-powered Autonomous Agents in Text-to-SQL tasks. MAC-SQL(Wang et al., 2023b) proposed an Agents framework for text2sql tasks, but it is a pipeline and non-autonomous Agents system, which does not have the flexibility and explorability of Autonomous Agents.
6.3. Complex Schema
In real-world Text-to-SQL tasks, complex table schemas are involved. An example contemplated by (Floratou and Psallidas, 2024) is where one of Microsoft’s internal financial data warehouses encompasses 632 tables with over 4000 columns and 200 views with more than 7400 columns. This introduces a torrent of tokens in the Text2SQL task, generating two prominent issues. The first issue lies in the difficulty of selecting correct tables and columns when generating SQLs since the vast context length leads to the diversion of LLM’s attention, hence resulting in ineffective SQL generation. On the other hand, the vast number of tokens prolongs the inference time of LLM. One possible solution to mitigate these issues is schema linking, as suggested in (Floratou and Psallidas, 2024). It states, better accuracy can be achieved by selecting only the relevant tables and columns related to the Natural Language (NL) query.
6.4. Benchmark
(Floratou and Psallidas, 2024) highlights that real-world Text-to-SQL tasks always have complex table schemas. However, commonly utilized standard datasets like Spider(Yu et al., 2018), among others, only contain simple table schemas, even though they have played an important role in driving the evolution of the text-to-SQL field. The Bird(Li et al., 2024a) dataset emulates real scenarios: multiple tables, complicated questions, including noise. However, when compared to the scale of tables in Microsoft’s internal financial data warehouses mentioned above, it still relatively simple.
6.5. Domain Knowledge
The training process of the Large Langeuage Model(LLM) involves a large amount of corpus information, which gives the LLM rich general knowledge, laying the foundation for LLM’s powerful text-to-SQL task capabilities. However, in the industry, text-to-SQL also requires the model to have task-related domain knowledge, such as the meaning of industry-specific jargon. LLM may have serious bias in understanding and answering questions without related domain knowledge.
LLM can acquire domain knowledge through two methods: prompt engineering and fine-tuning. The former is through Retrieval-Augmented Generation(RAG), while the latter uses domain knowledge-related training data to fine-tune the LLM. However, both methods have their own issues. For prompt engineering, we face two problems. One is in building the knowledge base. In reality, domain knowledge generally exists in documents, which are often unstructured and noisy, making it difficult to construct a structured and high-quality knowledge base for RAG’s usage. The other problem is in knowledge retrieval, where the commonly used similarity-based retrieval is not accurate enough and might include irrelevant information. Irrelevant information may degrade the performance of the LLM. For the fine-tuning method, there are mainly three issues. First, the training cost of fine-tuning is high. Second, fine-tuning could lead to catastrophic forgetting in the LLM. Third, the knowledge trained into the LLM is hard to modify because simply re-training the whole model on an updated set is expensive, while finetuning the model only with new data may affect a large number of other implicitly memorized facts.
7. Conclusion
This paper offers an comprehensive review of the utilization of Large Language Models (LLMs) in Text-to-SQL tasks. In this survey, the application of LLMs to Text-to-SQL tasks is categorized into two distinct categories: prompt engineering and finetuning. In the prompt engineering categoriy, we provide a detailed discussion revolving around the basic structure of prompts in Text-to-SQL tasks, the method of supplemental knowledge integration, selection of examples, and the reasoning process. In the finetuning categoriy, we begin with the basic steps of finetuning: data preparation, pre-training model selection, model finetuning, and model evaluation, each discussed specifically with Text-to-SQL tasks. In addition, we have reviewed the existing benchmarks for Text-to-SQL tasks and discuss the benchmark datasets suitable for the era of large Language models in detail. Lastly, we take a look at potential future directions while sharing our perspectives. We hope this survey will enable readers to gain a broader understanding of the recent advances in this field and offer some insights into its future trajectory.